(a) Find the probability that in a family of 5 children there will be (i) at least one boy, (ii) at least one boy and at least one girl. Assume that the probability of a male birth is 1/2. (b) Suppose a random variable X has the probability density f(x) =    x, 0 ≤ x < 1 2 − x, 1 ≤ x < 2 0, elsewhere Find (i) P(−1 < x < 0.5) (ii)P(x ≤ 1.5) (iii) P(x > 2.5) (iv) P(0.25 < x < 1.5)
1
Expert's answer
2021-10-26T11:28:46-0400

a) Let "X=" the number of boys in a family of 5 children: "x\sim Bin(n, p)"  

Given "n=5, p=\dfrac{1}{2}, q=1-\dfrac{1}{2}=\dfrac{1}{2}"


"P(at\ least\ 1\ boy)=1-P(X=0)"

"=1-\dbinom{5}{0}(\dfrac{1}{2})^5(\dfrac{1}{2})^{5-5}=\dfrac{31}{32}"

"P(at\ least\ 1\ boy\ and \ at\ least\ 1\ girl)"

"=P(X=1)+P(X=2)+P(X=3)+P(X=4)"

"=1-P(X=0)-P(X=5)"

"=1-\dbinom{5}{0}(\dfrac{1}{2})^5(\dfrac{1}{2})^{5-5}-\dbinom{5}{5}(\dfrac{1}{2})^0(\dfrac{1}{2})^{5-0}"

"=\dfrac{30}{32}=\dfrac{15}{16}"

b)


"f(x) = \begin{cases}
 1-x, &0\leq x<1 \\
 2-x, &1\leq x<2 \\
0, & elsewhere
\end{cases}"

(i)


"P(−1 < x < 0.5)=\displaystyle\int_{-1}^{0.5}f(x)dx=\displaystyle\int_{0}^{0.5}(1-x)dx"

"=[x-x^2/2]\begin{matrix}
 0.5 \\
 0
\end{matrix}=0.5-0.125=0.375"

(ii)


"P(x\leq1.5)=\displaystyle\int_{-\infin}^{1.5}f(x)dx"

"=\displaystyle\int_{0}^{1}(1-x)dx+\displaystyle\int_{1}^{1.5}(2-x)dx"

"=[x-x^2/2]\begin{matrix}
 1 \\
 0
\end{matrix}+[2x-x^2/2]\begin{matrix}
 1.5 \\
 1
\end{matrix}"

"=1-0.5+3-1.125-2+0.5=0.875"

(iii)


"P(x>2.5)=1"

(iv)


"P(0.25 < x < 1.5)=\displaystyle\int_{0.25}^{1.5}f(x)dx"

"=\displaystyle\int_{0.25}^{1}(1-x)dx+\displaystyle\int_{1}^{1.5}(2-x)dx"

"=[x-x^2/2]\begin{matrix}
 1 \\
 0.25
\end{matrix}+[2x-x^2/2]\begin{matrix}
 1.5 \\
 1
\end{matrix}"




"=1-0.5-0.25+0.03125+3-1.125-2+0.5"

"=0.65625"